# CUBIC DIOPHANTINE EQUATION SATISFIABILITY IS $\Sigma^{0}_{1}$-COMPLETE

### v6

@@ 6.3 Non-Enumerability of the Region
Theorem 6.5 (Region Impredicativity).
@@
$\mathcal{U}={(d,n)\mid\mathrm{H}{\mathbb{N}}(d,n)=\mathsf{U}}.$ (6.134)
@@
Then, (i) the Membership $(d,n)\in\mathcal{U}$ is a $\Pi^{0}_{2}$ predicate, (ii) for each $d$ there exists a minimal threshold
$n^{*}(d)=\min{n\mid(d,n)\in\mathcal{U}}\in\mathbb{N}\cup{\infty},$ (6.135)
@@
By Reductio ad absurdum in steps. We consult Kleene [1952],
@@
    $\forall e;\neg\Big(\mathrm{Tot}*{d,n}(e)\wedge\forall x,\big(\Phi_{e}(x)=1\leftrightarrow\mathrm{Sol}*{d,n}(x)\big)\Big),$ (6.136)
    on instances of type $(d,n)$. Totality is $\Pi^{0}_{2}$, and correctness is arithmetic; hence the whole statement is $\Pi^{0}_{2}$ in $(d,n)$. This proves (i).
@@
    For fixed $d$, if $\mathrm{H}{\mathbb{N}}(d,n)$ is undecidable, then so is $\mathrm{H}{\mathbb{N}}(d,n^{\prime})$ for all $n^{\prime}\geq n$. Hence each degree-slice of $\mathcal{U}$ is upward closed in (n). Therefore the minimal threshold $n^{*}(d)$ is well-defined in $\mathbb{N}\cup{\infty}$.
@@
    Suppose toward contradiction that $n^{*}$ is computable, observe how the predicate $(d,n)\in\mathcal{U}\Longleftrightarrow n\geq n^{*}(d)$ becomes decidable. Hence the entire undecidability region would admit a computable boundary, in that $d\longmapsto n^{*}(d)$ encodes the minimal variable count at which equations admit arbitrary $\Sigma^{0}_{1}$ predicates. Determining $n^{*}(d)$ would admit whether a fixed universal simulation can be compressed to $(d,n)$. Contradiction. More concretely: from an index $e$, one could effectively construct a degree $d_{e}$ encoding of $\Phi_{e}$ whose minimal variable count reflects whether $\Phi_{e}$ is total. If $n^{*}$ were computable, totality of arbitrary machines would be decidable. Hence $n^{*}$ cannot be computable. This establishes (ii), (iii) follows immediately.
Intuition.
@@
This has a structural consequence. Any formal specification of the “first” degree $k$ at which undecidability arises, if carried out inside a theory strong enough to internalize the reduction of Theorem 1.8, would interact nontrivially with the same reflection mechanism. The act of precisely defining the threshold is, by the reflection route of Theorem 5.6 and Corollary 5.7, the act of crossing it. Whether $\mathcal{U}$ is itself recursively enumerable remains ope. What the present construction does establish is that the boundary resists specification from within any theory strong enough to formalize the reduction. The question,
@@
is by the constructions shown, an instance of the very phenomenon it tries to pin down. The self-referential structure of a cubic closure principle suggests a stronger phenomenon. The reduction of Section 5 encodes itself at degree $3$: the very mechanism that witnesses $\Sigma^{0}_{1}$-completeness—the guard gadget, the proof-to-polynomial compiler, the diagonal fixpoint—is itself a cubic construction. One cannot “see” the cubic threshold without building the cubic encoding that establishes it. This motivates the following conjecture, understood less as a claim than as a structural bet, the act of precisely locating the undecidability boundary already require crossing it. We close with a bet that we cannot win, only lose by falsification.
Conjecture 6.6 (Diophantine Gap Conjecture).
@@
We suspect that $\operatorname{T}$ will not emit a $\rm D$} nor $\rm U$} for $\varphi$ in some arithmetized sentence expressing $\mathrm{H}{\mathbb{N}}(3,2)=\varphi$.

### v7

@@ 7.2 Non-Enumerability of the Region
Theorem 7.2.1 (Region Impredicativity).
@@
$\mathcal{U}=\{(d,n)\mid\mathrm{H}{\mathbb{N}}(d,n)=\mathsf{U}\}.$ (167)
@@
Then, (i) the Membership $(d,n)\in\mathcal{U}$ is a $\Pi^{0}_{\geq 2}$ predicate, (ii) for each $d$ there exists a minimal threshold
$n^{*}(d)=\min\{n\mid(d,n)\in\mathcal{U}\}\in\mathbb{N}\cup\{\infty\},$ (168)
@@
By reductio in steps. We consult Kleene [1952].
@@
    $\forall e\;\neg\Big(\mathrm{Tot}_{d,n}(e)\wedge\forall x\,\big(\Phi_{e}(x)=1\leftrightarrow\mathrm{Sol}_{d,n}(x)\big)\Big),$ (169)
    on instances of type $(d,n)$. Totality is $\Pi^{0}_{\geq 2}$, and correctness is arithmetic; hence the whole statement is $\Pi^{0}_{\geq 2}$ in $(d,n)$. This proves (i).
@@
    For fixed $d$, if $\mathrm{H}{\mathbb{N}}(d,n)$ is undecidable, then so is $\mathrm{H}{\mathbb{N}}(d,n^{\prime})$ for all $n^{\prime}\geq n$. Hence each degree-slice of $\mathcal{U}$ is upward closed in $n$. Therefore the minimal threshold $n^{*}(d)$ is well-defined in $\mathbb{N}\cup\{\infty\}$.
@@
    Suppose toward contradiction that $n^{*}$ is computable, observe how the predicate $(d,n)\in\mathcal{U}\Longleftrightarrow n\geq n^{*}(d)$ becomes decidable. Hence the entire undecidability region would admit a computable boundary, in that $d\longmapsto n^{*}(d)$ encodes the minimal variable count at which equations admit arbitrary $\Sigma^{0}_{1}$ predicates. Determining $n^{*}(d)$ would admit inconsistency. More concretely: from an index $e$, one could effectively construct a degree $d_{e}$ encoding of $\Phi_{e}$ whose minimal variable count reflects whether $\Phi_{e}$ is total. If $n^{*}$ were computable, totality of arbitrary machines would be decidable. Hence $n^{*}$ cannot be computable. This establishes (ii), (iii) follows immediately.
Remark 7.2.2.
@@
What the construction shows is that the boundary resists specification from within any theory strong enough to formalize the reduction. The question,
@@
is by the constructions shown, an instance of the very phenomenon it tries to pin down. The self-referential structure of a cubic closure principle suggests a stronger phenomenon. The reduction of Section 5 encodes itself at degree $3$: the very mechanism that witnesses $\Sigma^{0}_{1}$-completeness—the guard gadget, the proof-to-polynomial compiler, the diagonal fixpoint—is itself a cubic construction. One cannot “see” the cubic threshold without building the cubic encoding that establishes it. This motivates the following conjecture, understood less as a claim than as a structural bet, the act of precisely locating the undecidability boundary already requires crossing it. We close with a bet that we cannot win, only lose by falsification.
Conjecture 7.2.3 (Diophantine Gap Conjecture).
@@
We suspect that $\operatorname{T}$ will prove neither a $\rm D$} nor a $\rm U$} for $\varphi$ in some arithmetized sentence expressing $\mathrm{H}{\mathbb{N}}(3,2)=\varphi$.